OpenStudy (grant330sims):

Find the Dot Product: u=<3,2>, v=<-2,1>

OpenStudy (anonymous):

\[u=<3,2>\] \[v=<-2,1>\] \[u\cdot v=3*(-2)+2*(1)\] \[\boxed{\boxed{u\cdot v=-4}}\]

OpenStudy (grant330sims):

Thanks Dex! lol

OpenStudy (anonymous):

Nop ;D